OpenStudy (anonymous):

using the addition method,solve each system of equations 2x+9y=30 2x+y=-2

OpenStudy (anonymous):

again, lets multiply 2nd equation by -1

OpenStudy (anonymous):

2x +9y = 30 -2x -y = 2 8y = 32 y = 4 2x +y = -2 2x + 4 = -2 2x = -2 -4 2x = -6 x = -3 (-3, 4)

OpenStudy (anonymous):

wait wait where did u get the -2?

OpenStudy (anonymous):

nevermind i got that one :)

OpenStudy (anonymous):

K

OpenStudy (anonymous):

i need help with a few more.. i going to try to do it by myself but u can go ahead and do the same as u did for this one and ill re check myself :)